OpenStudy (anonymous):

solve using quadrtic formulia 3x^2 + 7x - 20 = 0

OpenStudy (helder_edwin):

for the equation \(ax^2+bx+c=0\) the quadratic formula is \[ \large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \]

OpenStudy (anonymous):

I think this one is factorable too...